OpenStudy (anonymous):

Is the sequence geometric? If so, identify the common ratio. 2, –4, –16, –36, ... yes; –2 yes; 2 yes; –3 no

OpenStudy (zzr0ck3r):

-4/2=2 -16/-4=4 so ?

OpenStudy (zzr0ck3r):

if \(\{a_0,a_1,a_2,....,a_{n-1},a_n,....\}\) is geometric, then \(\frac{a_1}{a_0}=\frac{a_2}{a_1}=......=\frac{a_n}{a_{n-1}}=.....\)

OpenStudy (anonymous):

Not geometric since the common ration changes? (Sorry for replying late the site wouldn't let me get on)

OpenStudy (zzr0ck3r):

correct